# Unlocking State-Tracking in Linear RNNs Through Negative Eigenvalues

### v1

@@ 4.3 Expressivity of Products of Generalized Householder Matrices
We first observe that if ${\bm{M}}\in\mathcal{M}_{1}^{n}(\{-1\})$, then ${\bm{M}}$ is a reflection (or Householder) matrix, and that for any ${\bm{x}}\in\mathbb{R}^{l}$, ${\bm{A}}_{\mathrm{GH}}({\bm{x}})\in\mathcal{M}_{1}^{n}([0,1])$ and ${\bm{A}}^{-}_{\mathrm{GH}}({\bm{x}})\in\mathcal{M}_{1}^{n}([-1,1])$ so that with our change we also include reflection matrices. Moreover, $\mathcal{M}_{k}^{n}(\Omega)\subseteq\mathcal{M}_{k^{\prime}}^{n}(\Omega^{\prime})$ if $1\in\Omega$, $k^{\prime}\geq k$ and $\Omega\subseteq\Omega^{\prime}$.
Our next result shows that products of GH matrices can represent any matrix with Euclidean norm less than or equal to 1. However, if every GH matrix in the product has only positive eigenvalues, matrices with complex eigenvalues cannot be represented. In contrast, repeated products of triangular matrices with eigenvalues in $[-1,1]$ remain triangular, with eigenvalues in the same range.
@@
    Any eigenvalue $\lambda$ of ${\bm{N}}\in\mathcal{M}_{k}^{n}((-1,1])$ is either $1$ or satisfies $|\lambda|<1$ and if in addition ${\bm{N}}\in\mathcal{M}_{k}^{n}([0,1])$, then $\lambda\in\mathbb{R}$.
@@
A consequence of Proposition 1 is that if for every layer of an LRNN, there exists $n,k\in\mathbb{N}$ such that the map ${\bm{A}}$ from inputs to state-transition matrices is such that ${\bm{A}}:\mathbb{R}^{l}\to\mathcal{M}_{k}^{n}((-1,1])$, then the LRNN cannot learn to count modulo $m$, with $m$ not a power of two, due to Theorem 2. In contrast, if we allow ${\bm{A}}:\mathbb{R}^{l}\to\mathcal{M}_{k}^{n}([-1,1])$ and $k$ is large enough, the following theorem shows that an LRNN with one layer can implement any FSA whose transition monoid is a group.
@@
LRNNs with state transition matrices that are repeated products of GH matrices, each with eigenvalues in the range $[-1,1]$, can recognize any regular languages. In particular, every FSA $\mathcal{A}=(\Sigma,Q,q_{0},\delta)$ can be implemented by a finite precision LRNN with $s\leq 2^{|Q|}$ layers, each of the form 1, where $n\leq|Q|$, $p\leq s$, $d=1$, ${\bm{A}}:\mathbb{R}^{l}\to\mathcal{M}_{n}^{n}([-1,1])$ and ${\bm{B}}:\mathbb{R}^{l}\to\mathbb{N}^{n}$.
The proof in Section C.4 exploits the landmark Theorem by Krohn & Rhodes (1965), which states that every FSA can be decomposed as a cascade of simpler FSAs whose state-transition functions are either one-to-one or constant. Each layer of the LRNN will implement one FSA (with $n$ states) of the cascade using $n\times n$ permutation matrices, which are in $\mathcal{M}_{n-1}^{n}(\{-1,1\})$, for the one-to-one transitions, while for constant (state-independent) transitions it will set the corresponding state-transition matrix to $0\in\mathcal{M}_{n}^{n}(\{0\})$ and the function ${\bm{B}}$ appropriately.
Note, that we can obtain the zero matrix only inefficiently as a product of $n$ GH matrices, while it could also be obtained with a single diagonal matrix. This points towards hybrids LRNNs using a mix of GH and diagonal matrices, whose exploration we leave for future work.
@@ 5 Experiments
Mamba $\mathrm{Diag}({\bm{s}}({\bm{x}}_{t}))$ $\mathrm{Diag}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2}{\bm{s}}({\bm{x}}_{t})-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1})$
@@ 5.2 State-Tracking
We perform experiments on group word problems, relying on the code provided by Merrill et al., 2024. In particular, we focus on the $S_{5}$ group, which is the first unsolvable symmetric group where current LRNN and Transformers are known to perform poorly. We also report results for the addition modulo $60$, i.e. the cyclic group $\mathbb{Z}_{60}$, in Section E.2.2. We note that parity is $S_{2}$. In these experiments, the input to the model is a sequence of group elements, while the supervision is given by another sequence of group elements, each being the product of the previous ones in the input. Since solving $S_{5}$ would require LRNNs with state-transition matrices that are repeated products of 4 GH matrices (see Theorem 3), each with eigenvalues $[-1,1]$, we also consider three simplified setups: (i) allowing as inputs only permutations up to 2 elements (identity and swaps), (ii) allowing only permutations up to 3 elements, (iii) using 4 tokens for each permutation. Additional details are in Section E.2. We stress that, even when restricting the inputs, possible outputs remain the same, since swaps are generators of the group.
@@ 5.3 Language Modeling
Results  All models trained stably with our modification and without changing the learning rate. The validation perplexity of the proposed variants was comparable, albeit slightly worse than that of the original models throughout training (see Figure 7 in the Appendix). The experiments in Figure 3 demonstrate that on coding and math datasets, DeltaNet with an eigenvalue range of $[-1,1]$ achieves lower perplexity than the baseline with range $[0,1]$ for both model sizes. For TriviaQA, the perplexity of DeltaNet $[-1,1]$ is slightly higher. Note, that this is a task relying on memorization, not linked to state-tracking, and hence we do not expect an improvement. On SlimPajama, we also observe slight improvement with our modification. For Mamba instead, our modifications consistently degrades the performance on these tasks (Figure 8 in the Appendix). To ensure that our models are comparable with those obtained by Yang et al. (2024b), we evaluate them on the same benchmark tasks from lm-harness (Gao et al., 2024) in Table 4. Note, that we trained on 100B tokens of FineWeb, while Yang et al. (2024b) reported results from training on 15B and 100B tokens of SlimPajama.
At 340-370M parameters, with the extended range both architectures show enhanced performance in some of the tasks: Mamba in the second subset of tasks (+2.1% average accuracy) and DeltaNet in retrieval tasks (+2% SWDE, +4.4% SQUAD). At 1.3B parameters, extending the eigenvalue range of DeltaNet shows mixed results, suggesting that the increased expressivity may need training beyond 100B tokens to fully unlock the model’s capacity.
@@ 6 Conclusion
In this work, we showed the substantial impact of extending the eigenvalue range of state-transition matrices in LRNNs from $[0,1]$ to $[-1,1]$. This modification provably enhances LRNN expressivity in state-tracking tasks, without adding overhead in training or inference. While Mamba successfully solves the parity problem, its diagonal matrix structure limits further performance gains. In contrast, DeltaNet, by leveraging its non-diagonal matrix structure enabling simultaneous token and channel mixing, excels across a broader spectrum of tasks. Our results underscore the critical role of non-diagonal state-transition matrices in augmenting state-tracking capabilities, highlighting a promising direction for future LRNN advancements.
Limitations and Future work  Our modification is not directly compatible with a numerical technique used by some diagonal LRNNs such as Mamba2, GLA and mLSTM. In partticular, these models rely on positive state-transition matrices to compute cumulative products in log space, which improves numerical accuracy and potentially training stability (see Section E.4 for details). Further research is needed to assess the impact of training large-scale language models with state-tracking capabilities. To this end, we aim to understand the potential downsides of increased expressivity. For example, we hypothesize a fundamental trade-off between state-tracking and associative recall, which is also of theoretical interest and could guide hybrid model design. Moreover, the theoretical expressivity of DeltaNet $[-1,1]$ with multiple layers is still unclear. We showed that it can solve addition modulo $m$ (in Appendix D) which is equivalent to the $\mathbb{Z}_{3}$ group word problem, but we do not know if it can also solve word problems for the symmetric groups $\mathbb{S}_{n}$ with $n\geq 3$.
@@ Acknowledgments
We would like to thank David Salinas, Eric Alcaide, Arya Akhavan, Matia Bojovic, and the active members of the Flash Linear Attention discord channel for their constructive discussions and feedback. This research was partially supported by the following sources: PNRR MUR Project PE000013 CUP J53C22003010006 “Future Artificial Intelligence Research (FAIR)“, funded by the European Union – NextGenerationEU, and EU Project ELSA under grant agreement No. 101070617. TAILOR, a project funded by EU Horizon 2020 research and innovation programme under GA No 952215; the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under grant number 417962828; the European Research Council (ERC) Consolidator Grant “Deep Learning 2.0” (grant no. 101045765). Frank Hutter acknowledges financial support by the Hector Foundation. The authors acknowledge support from ELLIS and ELIZA. Funded by the European Union. The authors gratefully acknowledge the Gauss Center for Supercomputing eV (www.gauss-centre.eu) for funding this project by providing computing time on the GCS supercomputer JUWELS at Jülich Supercomputing Center (JSC). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the ERC. Neither the European Union nor the ERC can be held responsible for them.
@@
  • Appendix A contains additional additional details on the notation used, the relationship between RNNs and regular languages, the assumption of finite precision, the states, and the decoder function.
@@ A.1 Notation
We denote with $\mathbb{C},\mathbb{R},\mathbb{N}$ the sets of complex, real, and natural numbers, respectively. We use lowercase letters for scalar quantities (e.g. $x\in\mathbb{R}$), bold lowercase letters for (column) vectors (e.g. ${\bm{v}}\in\mathbb{R}^{n}$), and bold uppercase letters for matrices (e.g. ${\bm{M}}\in\mathbb{R}^{n\times d}$). Some functions with matrix (vector) outputs, such as ${\bm{A}}$ and ${\bm{B}}$ in (1), are also bold upper (lower) case letters to emphasize the fact that they output matrices (vectors). We denote with $\left\lVert{\bm{v}}\right\rVert$ the Euclidean norm of the vector ${\bm{v}}\in\mathbb{R}^{n}$. When ${\bm{M}}\in\mathbb{R}^{n\times d}$, $\left\lVert{\bm{M}}\right\rVert$ also refers to the Euclidean norm, corresponding to the largest singular value. The vector ${\bm{e}}_{i}\in\mathbb{R}^{n}$ is the $i$-th vector of the canonical bases in $\mathbb{R}^{n}$, i.e. the one-hot vector with $1$ only in the $i$-th component and 00 in the others.
We also define for a Boolean $s$
$\mathbf{1}\{s\}:=\begin{cases}1\text{ if $s$ is true}\\ 0\text{ if $s$ is false }.\end{cases}$
@@
A.2 Regular Languages and Recurrent Neural Networks

### v2

@@ 4.3 Expressivity of Products of Generalized Householder Matrices
We observe that if ${\bm{M}}\in\mathcal{M}_{1}^{n}(\{-1\})$, then ${\bm{M}}$ is a reflection (or Householder) matrix, and that for any ${\bm{x}}\in\mathbb{R}^{l}$, ${\bm{A}}_{\mathrm{GH}}({\bm{x}})\in\mathcal{M}_{1}^{n}([0,1])$ and ${\bm{A}}^{-}_{\mathrm{GH}}({\bm{x}})\in\mathcal{M}_{1}^{n}([-1,1])$ so that with our change we also include reflections. Moreover, $\mathcal{M}_{k}^{n}(\Omega)\subseteq\mathcal{M}_{k^{\prime}}^{n}(\Omega^{\prime})$ if $\Omega\subseteq\Omega^{\prime}$ and either $k^{\prime}=k$ or $k^{\prime}\geq k$, $1\in\Omega$.
Our next result shows that products of GH matrices can represent any matrix with Euclidean norm less than or equal to 1. However, there are some restriction if we only allow each GH matrix in the product to have positive eigenvalues. In contrast, repeated products of upper (lower) triangular matrices with eigenvalues in $[-1,1]$ remain upper (lower) triangular, with eigenvalues in the same range.
@@
    Any eigenvalue $\lambda$ of ${\bm{N}}\in\mathcal{M}_{k}^{n}((-1,1])$ is either $1$ or satisfies $|\lambda|<1$ and if in addition ${\bm{N}}\in\mathcal{M}_{k}^{n}([0,1])$ and $k\leq 2$, then $\lambda\in[0,1]\subset\mathbb{R}$.
@@
A consequence of Proposition 1 is that if for every layer of an LRNN, there exists $n,k\in\mathbb{N}$ such that the map ${\bm{A}}$ from inputs to state-transition matrices is such that ${\bm{A}}:\mathbb{R}^{l}\to\mathcal{M}_{k}^{n}([0,1])$, then if $k\leq 2$ the LRNN cannot learn parity, due to Theorem 1, and for any $k$ the state transition matrices are either the identity or not orthogonal, and hence unable to represent reflections or rotations. In contrast, if we allow ${\bm{A}}:\mathbb{R}^{l}\to\mathcal{M}_{k}^{n}([-1,1])$ and $k$ is large enough, the following theorem shows that an LRNN with one layer can implement any FSA whose transition monoid is a group, and that $n=k=2$ is enough for cyclic groups (modular addition).
@@
LRNNs with state transition matrices that are repeated products of GH matrices, each with eigenvalues in the range $[-1,1]$, can recognize any regular language. In particular, every FSA $\mathcal{A}=(\Sigma,Q,q_{0},\delta)$ can be implemented by a finite precision LRNN with $s\leq 2^{|Q|}$ layers, each of the form 1, where $n\leq|Q|$, $p\leq s$, $d=1$, ${\bm{A}}:\mathbb{R}^{l}\to\mathcal{M}_{n}^{n}([-1,1])$ and ${\bm{B}}:\mathbb{R}^{l}\to\mathbb{N}^{n}$.
The proof in Section C.4 exploits the landmark Theorem by Krohn & Rhodes (1965), which states that every FSA can be decomposed as a cascade of simpler FSAs whose state-transition functions are either one-to-one or constant. Each layer of the LRNN will implement one FSA (with $n$ states) of the cascade using $n\times n$ permutation matrices, which are in $\mathcal{M}_{n-1}^{n}(\{-1,1\})$, for the one-to-one transitions, while for constant (state-independent) transitions it will set the corresponding state-transition matrix to $0\in\mathcal{M}_{n}^{n}(\{0\})$ and the function ${\bm{B}}$ appropriately. Note, that we can obtain the zero matrix only inefficiently as a product of $n$ GH matrices, while it could also be obtained with a single diagonal matrix. This points towards hybrids LRNNs using a mix of GH and diagonal matrices, whose exploration we leave for future work.
@@ 5 Experiments
Mamba $\mathrm{Diag}({\bm{s}}({\bm{x}}_{t}))$ $\mathrm{Diag}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2}{\bm{s}}({\bm{x}}_{t}){\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}-1})$
@@ 5.2 State-Tracking
We perform experiments on group word problems, relying on the code provided by Merrill et al., 2024. https://github.com/jopetty/word-problem. In particular, we focus on the $S_{5}$ group, which is the first unsolvable symmetric group where current LRNN and Transformers are known to perform poorly. We also report results for the addition modulo $60$, i.e. the cyclic group $\mathbb{Z}_{60}$, in Section E.2.2. We note that parity is $S_{2}$. In these experiments, the input to the model is a sequence of group elements, while the supervision is given by another sequence of group elements, each being the product of the previous ones in the input. Since solving $S_{5}$ would require LRNNs with state-transition matrices that are repeated products of 4 GH matrices (see Theorem 3), each with eigenvalues $[-1,1]$, we also consider three simplified setups: (i) allowing as inputs only permutations up to 2 elements (identity and swaps), (ii) allowing only permutations up to 3 elements, (iii) using 4 tokens for each permutation. Additional details are in Section E.2. We stress that, even when restricting the inputs, possible outputs remain the same, since swaps are generators of the group.
@@ 5.3 Language Modeling
Results  All models trained stably with our modification and without changing the learning rate. The validation perplexity of the proposed variants was comparable, albeit slightly worse than that of the original models throughout training (see Figure 7 in the Appendix). The experiments in Figure 3 demonstrate that on coding and math datasets, DeltaNet with an eigenvalue range of $[-1,1]$ achieves lower perplexity than the baseline with range $[0,1]$ for both model sizes. For TriviaQA, the perplexity of DeltaNet $[-1,1]$ is slightly higher. Note, that this is a task relying on memorization, not linked to state-tracking, and hence we do not expect an improvement. On SlimPajama, we also observe slight improvement with our modification. For Mamba instead, our modifications consistently degrades the performance on these tasks (Figure 8 in the Appendix).
To ensure that our models are comparable with those obtained by Yang et al. (2024b), we evaluate them on the same benchmark tasks from lm-harness (Gao et al., 2024) in Table 4. Note, that we trained on 100B tokens of FineWeb, while Yang et al. (2024b) reported results from training on 15B and 100B tokens of SlimPajama. At 340-370M parameters, with the extended range both architectures show enhanced performance in some of the tasks: Mamba in the second subset of tasks (+2.1% average accuracy) and DeltaNet in retrieval tasks (+2% SWDE, +4.4% SQUAD). At 1.3B parameters, extending the eigenvalue range of DeltaNet shows mixed results, suggesting that the increased expressivity may need training beyond 100B tokens to fully unlock the model’s capacity.
@@ 6 Conclusion
In this work, we showed the substantial impact of extending the eigenvalue range of state-transition matrices in LRNNs from $[0,1]$ to $[-1,1]$. This modification provably enhances LRNN expressivity in state-tracking tasks, without adding overhead in training or inference. While Mamba successfully solves the parity problem, its diagonal matrix structure limits further gains. In contrast, DeltaNet, thanks to its non-diagonal state transition matrices which enable simultaneous token and channel mixing, excels across a broader spectrum of tasks. Our results underscore the critical role of non-diagonal state-transition matrices in augmenting state-tracking capabilities, highlighting a promising direction for future LRNN advancements.
Limitations and Future work  Our modification is not directly compatible with a numerical technique used by some diagonal LRNNs such as Mamba2, GLA and mLSTM. In particular, these models rely on positive state-transition matrices to compute cumulative products in log space, which improves numerical accuracy and potentially training stability (see Section E.4 for details). Further research is needed to assess the impact of training large-scale language models with state-tracking capabilities. To this end, we aim to understand the potential downsides of increased expressivity. For example, we hypothesize a fundamental trade-off between state-tracking and associative recall, which is also of theoretical interest and could guide hybrid model design. Moreover, the theoretical expressivity of DeltaNet $[-1,1]$ with multiple layers is still unclear. We showed that it can solve addition modulo $m$ (in Appendix D) which is equivalent to the $\mathbb{Z}_{3}$ group word problem, but we do not know if it can also solve word problems for the symmetric groups $\mathbb{S}_{n}$ with $n\geq 3$.
@@ Acknowledgments
We would like to thank David Salinas, Herilalaina Rakotoarison, Eric Alcaide, Arya Akhavan, Matia Bojovic, Erfan Mirzaei and the active members of the Flash Linear Attention discord channel for their constructive discussions and feedback. We also wish to thank Mattia Pini for helping us with the Franklin computing cluster at IIT, on which we run our largest experiments. This research was partially supported by the following sources: PNRR MUR Project PE000013 CUP J53C22003010006 “Future Artificial Intelligence Research (FAIR)“, funded by the European Union – NextGenerationEU, and EU Project ELSA under grant agreement No. 101070617. TAILOR, a project funded by EU Horizon 2020 research and innovation programme under GA No 952215; the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under grant number 417962828; the European Research Council (ERC) Consolidator Grant “Deep Learning 2.0” (grant no. 101045765). Frank Hutter acknowledges financial support by the Hector Foundation. The authors acknowledge support from ELLIS and ELIZA. Funded by the European Union. The authors gratefully acknowledge the Gauss Center for Supercomputing eV (www.gauss-centre.eu) for funding this project by providing computing time on the GCS supercomputer JUWELS at Jülich Supercomputing Center (JSC). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the ERC. Neither the European Union nor the ERC can be held responsible for them.
@@
  • Appendix A contains additional details on the notation used, on Table 1, the relationship between RNNs and regular languages, the assumption of finite precision, the states, and the decoder function.
@@ A.1 Notation
We denote with $\mathbb{C},\mathbb{R},\mathbb{N}$ the sets of complex, real, and natural numbers, respectively. We use lowercase letters for scalar quantities (e.g. $x\in\mathbb{R}$), bold lowercase letters for (column) vectors (e.g. ${\bm{v}}\in\mathbb{R}^{n}$), and bold uppercase letters for matrices (e.g. ${\bm{M}}\in\mathbb{R}^{n\times d}$). Some functions with matrix (vector) outputs, such as ${\bm{A}}$ and ${\bm{B}}$ in (1), are also bold upper (lower) case letters to emphasize the fact that they output matrices (vectors). We use $\odot$ to indicate the element-wise (Hadamard) product between two vectors or matrices. We denote with $\left\lVert{\bm{v}}\right\rVert$ the Euclidean norm of the vector ${\bm{v}}\in\mathbb{R}^{n}$. When ${\bm{M}}\in\mathbb{R}^{n\times d}$, $\left\lVert{\bm{M}}\right\rVert$ also refers to the Euclidean norm, corresponding to the largest singular value. The vector ${\bm{e}}_{i}\in\mathbb{R}^{n}$ is the $i$-th vector of the canonical bases in $\mathbb{R}^{n}$, i.e. the one-hot vector with $1$ only in the $i$-th component and 00 in the others. We define the binomial coefficient for every $k,j\in\mathbb{N}$ with $j\leq k$ as
$\binom{k}{0}:=1,\quad\binom{k}{j}:=\frac{k(k-1)\dots(k-j+1)}{j!}.$
We also define for a Boolean $s$ and $x\in\mathbb{R}$
$\mathbf{1}\{s\}:=\begin{cases}1\text{ if $s$ is true}\\ 0\text{ if $s$ is false }\end{cases},\qquad\mathrm{sign}(x):=\begin{cases}1&\text{if }x\geq 0\\ -1&\text{if }x<0\end{cases}.$
@@
A.2 Details of Table 1
The Mamba recurrence in Equations 3 and 4 in (Gu & Dao, 2023) is applied independently to each channel of the input sequence. Expressing the full recurrence in the matrix-form of (1) is challenging, as it would require concatenating the rows of the matrix ${\bm{H}}_{t}$. For simplicity, in Table 1 we write instead the recurrence for each row of ${\bm{H}}_{t}$. In particular, Let ${\bm{x}}_{t}\in\mathbb{R}^{d}$ be the input of the layer, ${\bm{W}}_{\Delta}\in\mathbb{R}^{d\times d}$, ${\bm{w}}_{2}\in\mathbb{R}^{d}$, ${\bm{W}}_{1}=({\bm{w}}_{1},\dots,{\bm{w}}_{n})^{\top}\in\mathbb{R}^{n\times d}$ be learnable parameters, ${\bm{q}}_{t}\in\mathbb{R}^{n},{\bm{k}}_{t}=(k_{t,1},\dots,k_{t,n})^{\top}\in\mathbb{R}^{n}$ be learnable functions of the input and $\bm{\Delta}_{t}=\mathrm{softplus}({\bm{W}}_{\Delta}{\bm{x}}_{t})$. Then, if we set ${\bm{H}}_{t}=({\bm{h}}_{t,1},\dots,{\bm{h}}_{t,n})^{\top}\in\mathbb{R}^{n\times d}$ and ${\bm{H}}_{0}=0$, we can write the recurrence for the $i$-th row of ${\bm{H}}_{t}$ and the output as
${\bm{h}}_{t,i}={\bm{A}}_{i}({\bm{x}}_{t}){\bm{h}}_{t-1,i}+{\bm{B}}_{i}({\bm{x}}_{t}),\qquad\hat{\bm{y}}_{t}=\psi({\bm{H}}_{t}^{\top}{\bm{q}}_{t}+{\bm{w}}_{2}\odot{\bm{x}}_{t}))$
where ${\bm{A}}_{i}({\bm{x}}_{t})$ and ${\bm{B}}_{i}({\bm{x}}_{t})$ are the matrices stated in Table 1, i.e.
${\bm{A}}_{i}({\bm{x}}_{t}):=\mathrm{Diag}\left(\exp\left(-\bm{\Delta}_{t}\odot\exp({\bm{w}}_{1,i})\right)\right)\in\mathbb{R}^{d\times d},\qquad{\bm{B}}_{i}({\bm{x}}_{t}):=k_{t,i}\bm{\Delta}_{t}\odot{\bm{x}}_{t}\in\mathbb{R}^{d}.$
Alternatively, as done in (Yang et al., 2024b, Table 4), one could write the full matrix recurrence as:
${\bm{H}}_{t}=\underbrace{\exp\left(-\bm{\mathbf{1}\Delta}_{t}^{\top}\odot\exp({\bm{W}}_{1})\right)}_{{\bm{A}}({\bm{x}}_{t})}\odot{\bm{H}}_{t-1}+\underbrace{{\bm{k}}_{t}(\bm{\Delta}_{t}\odot{\bm{x}}_{t})^{\top}}_{{\bm{B}}({\bm{x}}_{t})}.$
where $\mathbf{1}$ is the vector of $n$ ones. However, such a recurrence is not in the form (1), since we have replaced the matrix-matrix product ${\bm{A}}({\bm{x}}_{t}){\bm{H}}_{t}$ with the element-wise product ${\bm{A}}({\bm{x}}_{t})\odot{\bm{H}}_{t}$. Note that we follow the implementation of ${\bm{B}}({\bm{x}}_{t})$ used in the official Mamba codebase, which simplifies the expression originally presented in Equation 4 of (Gu & Dao, 2023) as described by the authors in a GitHub Issue [^3] https://github.com/state-spaces/mamba/issues/19.
A.3 Regular Languages and Recurrent Neural Networks